(a) (–7) – 8 – (–25)
(b) (–13) + 32 – 8 – 1
(c) (–7) + (–8) + (–90)
 
 
smrbiombhrinietnbumtuibmsiodbm,wirjbhmriujmiosbm,

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Stdent
(a) (-7) – 8 – (-25)
= -7 -8 + 25
= -15 + 25
= 10

b) (-13) + 32 – 8 – 1
= -13 + 32 – 8 – 1
= 32 – 22
= 10

(c) (-7) + (-8) + (-90)
= -7 – 8 – 90
= - 105




Thanks